OpenStudy (anonymous):

Jude has seven times as many dimes as nickels. Their value is $5.25. How many nickels does he have? a. 49 b. 15 c. 7

OpenStudy (asnaseer):

call the number of nickels n then how many dimes does he have in terms of n?

OpenStudy (anonymous):

n/10?

OpenStudy (btaylor):

d = 7n 10d + 5n = 525 now, substitute in your value of d -->10(7n) + 5n = 525 70n + 5n = 525 75n = 525 n = 7 (and 49 dimes)

OpenStudy (anonymous):

i mean n/7

OpenStudy (asnaseer):

the questions says he has seven times as many dimes as nickels

OpenStudy (anonymous):

oh, stupid me read it wrong, i thought he has seven times as many nickles as dimes

OpenStudy (asnaseer):

:)

OpenStudy (anonymous):

i thought it was 49 but when i checked it i was like "what??" now i know why xD

OpenStudy (asnaseer):

thats good - at least you've learnt something :)

OpenStudy (anonymous):

thanks @btaylor

OpenStudy (anonymous):

you too @asnaseer

OpenStudy (asnaseer):

yw :)